soy bean meal is 12%protein corn meal is 6%protein. how many pounds of each should be mixed together in order to get 240 pounds mixture that is 8% protein
how many pounds of the cornmeal___________?
How many pounds of the soybean_______________?

Answer by checkley77(12844)

.12x+.06(240-x)=.08*240
.12x+14.4-.06x=19.2
.06x=19.2-14.4
.06x=4.8
x=4.8/.06
x=80 pounds of soy meal is needed.
240-80=160 pounds of corn meal is needed.
Proof:
.12*80+.06*160=.08*240
9.6+9.6=19.2
19.2=19.2